Find the two repeated limits of the function f(x,y)=(y-x/y+x)(1+x^2/1+y^2) at (0,0).Does the simultaneous limit of f exist as (x,y) -(0,0)?Give reasons for your answer
1
Expert's answer
2020-03-18T15:10:20-0400

The given function is

"f(x,y)=\frac{(y-x)}{(y+x)} \frac{(1+x^2)}{(1+y^2)}"

The repeated limits of "f(x,y)" are following:

"\lim x\to0(\lim y\to0 \space f(x,y) )=\lim x\to0(-1)(1+x^2)"

"=-1"

"\lim y\to 0(\lim x\to 0 \space f(x,y))=\lim y\to0 (\frac{1}{1+y^2})"

"=1"

Simultaneous limit of "f(x,y)" at "(0,0)" does not exist because repeated limits are not equal.